OpenStudy (anonymous):

@awesomeb You have 1200 feet of fencing and you want to make two fenced in enclosures by splitting one enclosure in half. What are the largest dimensions of this enclosure that you could build?

OpenStudy (anonymous):

Area= x(1200-2x)/3 Area= (-2x^2+1200x)/3 Area= 1/3(-2x^2+1200x) Thus, dA/dx= 1/3(-4x+1200) Maximum Area occurs when dA/dx=0 Setting 1/3(-4x+1200)=0, -4x+1200=0 (multiplied both sides by 3) -4x= -1200 4x=1200 divide by 4 x=300 The width will be (1200-2x) / 3 (1200-600) / 3 = 600/3 = 200 The largest dimension is 300 by 200 feet

OpenStudy (anonymous):

@awesomeb you're a total life saver :D Thank youuuuuuuuuuuuuuu

OpenStudy (anonymous):

np :)